OpenStudy (anonymous):

find the sum 10/15 + 4/14

OpenStudy (anonymous):

Start by finding the common denominator.

OpenStudy (anonymous):

and we add those

OpenStudy (anonymous):

or how do we do that

OpenStudy (anonymous):

You can only add fractions with a common denominator. The common denominator is the lowest number that can divide into both number evenly.

OpenStudy (anonymous):

I like the method of Upside-Down Division: \[\begin{arry}

OpenStudy (anonymous):

this is hard i dont get any of this

OpenStudy (anonymous):

At some point you will want to google upside down division. For now, I will tell you that these denominators have no common factor, therefore, the common factor is the product of both numbers. Multiply them together.

OpenStudy (anonymous):

Then, write equivalent fractions, with the new denominator. At that point, you can add them.

OpenStudy (anonymous):

wait you mean multiply the top or bottom

OpenStudy (anonymous):

Multiply the denominators (bottom numbers) to find the common denominator.

OpenStudy (anonymous):

ok and what do we do after that

OpenStudy (anonymous):

Then, do to the numerator, what you did to the denominator to get the equivalent fraction.

OpenStudy (anonymous):

ok so i got 210 for the bottom and i got to multiply the top too

OpenStudy (anonymous):

\(\frac{10}{15}+\frac{4}{14}\) Optional: reduce the fractions so you have smaller numbers to work with: \(\frac{2}{3} + \frac{2}{7}\) ...to be continued...

OpenStudy (anonymous):

see how did you get that ugh i dont get this :/

OpenStudy (anonymous):

Then, write equivalent fractions: \(\frac{2}{3} \times \frac{7}{7} + \frac{2}{7} \times \frac{3}{3}\)

OpenStudy (anonymous):

10 and 15 both divide by 5 so \(10 \div 5 = 2\) and \(15 \div 5 = 3\)

OpenStudy (anonymous):

oh ok and then how do we do the last one you sent me

OpenStudy (anonymous):

You can reduce a fraction by dividing the numerator and denominator by the same number (called a common factor.)

OpenStudy (anonymous):

You understand how to reduce fractions?

OpenStudy (anonymous):

uhh not really lol

OpenStudy (anonymous):

ur answer is 20/21

OpenStudy (anonymous):

\(\frac{2}{3} \times \frac{7}{7} + \frac{2}{7} \times \frac{3}{3}\) Multiplying a fraction (or any other number by one doesn't change it's value. Any number divided by itself equals 1. Same goes for division.

OpenStudy (anonymous):

@nickirivera How does giving the answer help a person to learn how to do it?

OpenStudy (anonymous):

umm. well maybe you could have given the answer first and then showed the person how to do it and work backwards? I'm sorry for interrupting but you did not have to report me...I didn't know that you were still trying to teach.. I thought the lesson was over.

OpenStudy (anonymous):

i like @nickirivera method better sorry

OpenStudy (anonymous):

@nickirivera I did not report you, just asked a question. @eddie16 Taking someone else's answer may be easier for tonight, but will cause you problems at test time. Your choice thought. Good luck.